State two conditions for the function f(x,y) to be continuous joint probability distribution of continuous random variables X and Y


1
Expert's answer
2021-10-11T10:07:32-0400

If random variables "X" and "Y" are continuous and are defined on the same sample space "S", then, their joint probability density function(pdf) is a continuous function given by,

"f(x,y)". The function "f(x,y)" must satisfy the following conditions.

  • "f(x,y)\geqslant0,\forall(x,y)\in\R^2"
  • "\displaystyle\iint_{\R^2}f(x,y)dxdy=1"


where "\R^2" is the set of pairs of real numbers given as,

"\R^2=\R*\R=\{(x,y):x\space and \space y\space are\space real\space numbers\}"